OpenStudy (anonymous):

you toss 2 coins. find P(at least 1 tail is tossed)

OpenStudy (anonymous):

P( no tail ) = 1/2*1/2=1/4 So P( at least 1 tail) is the opposite case of that

OpenStudy (anonymous):

(H,H) , (T,T) , (H,T) , (T.H) ==== POSSIBLE OUTCOMES

hartnn (hartnn):

1-P(none is tails)=1-(P(1st is head)and P(2nd is head))....=1-0.5*0.5

OpenStudy (anonymous):

so its 0.5?

hartnn (hartnn):

no....1-0.5*0.5=1-0.25=0.75 or 3/4

OpenStudy (anonymous):

2 coins possible outcomes are: Heads, Tails Tails, Heads Heads, Heads Tails, Tails Therefore, the probability of at least 1 tail from 2 coins is tossed is 3/4.